### v1

# Build An Influential Bot In Social Media Simulations With Large Language Models

Abstract
Understanding the dynamics of public opinion evolution on online social platforms is critical for analyzing influence mechanisms. Traditional approaches to influencer analysis are typically divided into qualitative assessments of personal attributes and quantitative evaluations of influence power. In this study, we introduce a novel simulated environment that combines Agent-Based Modeling (ABM) with Large Language Models (LLMs), enabling agents to generate posts, form opinions, and update follower networks. This simulation allows for more detailed observations of how opinion leaders emerge. Additionally, we present an innovative application of Reinforcement Learning (RL) to replicate the process of opinion leader formation. Our findings reveal that limiting the action space and incorporating self-observation are key factors for achieving stable opinion leader generation. The learning curves demonstrate the model’s capacity to identify optimal strategies and adapt to complex, unpredictable dynamics.
keywords:
Agent-Based Modelling, Large Language Model, Reinforcement Learning
[…]
1 Introduction
The emergence of online social media has revolutionized public opinion measurement and representation. Traditional macroscopic approaches to analyzing social phenomena often struggle to capture the intricacy, discontinuity and heterogeneity in individual behaviors.
Agent-Based Modeling(ABM) has proven effective in analyzing and explaining social phenomena in a more detailed level. ABMs apply experimental control over large population, enabling researchers to capture the nonlinear societal mechanisms and reveal large-scale societal emergenceJackson et al. (2017). Various rules within ABM have been developed to explain the phenomena of consensus, clustering and bi-polarity in public opinion.
However, ABMs primarily function through mathematical representations and do not inherently integrate natural language entities. Our research aim to bridge this gap between computational simulations and linguistic data by integrating ABM and Large Language Models(LLMs). LLMs have been used as a source of simulated respondents in various studies, demonstrating the ability to predict user’s future behaviour with a degree of accuracy.
In our research, we have developed a simulated environment that replicates the dynamics of group discussions on a specific topic in an online social media context. This environment is structured with agents representing individual participant and follow relationships as links. Agents release topic-related posts to represent the current opinion and the links will be updated due to opinion states. LLMs are used for post generation and opinion elicitation for agents. The generated posts are not only relevant to the topic but also reflect the complex influences present in a social media environment. The simulation result aligns with the polarity distribution of the actual dataset, indicating that our simulated environment reflects realistic discussion patterns.
Given the simulated environment that indicate the full opinion evolution process of a discussion on one topic. Now we move to investigate in individual behaviors of one particular target agent.
Reinforcement Learning(RL) is a simulation method which has the capability to mimic the human-like action selection. Using RL model, the target agent learn optimal strategies based on the interaction with environment. We specifically focus on training RL models on a target agent with the objective of maximizing its follower count in the simulated network. Our study explores various environment and observability settings to understand their impact on the agent’s ability to achieve this goal. In our designed cases , the convergence in RL reward learning curves indicating RL achieved optimal solutions.
The potential applications of this research are diverse across multiple domains. Marketing firms could leverage the model to predict the effectiveness of promotional strategies by simulating how opinions evolve in response to targeted campaigns. For instance, social acceptance of emerging technologies like 6G is crucial for business and government (Briguglio et al. (2021)), policymakers could use the tool to simulate and evaluate the public’s response to communication strategies. For cyber security, this research could assist in identifying malicious activities such as influence campaigns or the spread of misinformation.
Here we summarise the major contributions of this paper:
  • A novel application of RL to opinion leader generation in a simulated social media environment.
  • The findings that limiting the action space and incorporating self-observation are key factors for stable generation under varying environments.
2 Related Work
2.1 Opinion Leader
2.2 Agent-Based Modelling
Agent-based modelling has been widely used by social researchers in the field of social simulation. In the modelling, the agents represent human individuals, interact with each other and with the system environment explicitly and individually. Compared to the other tools, such as laboratory or filed experiments, ABMs apply the experimental control on large scale population to capture the nonlinear societal mechanisms and reveal large-scale societal emergence (Jackson et al. (2017)).
However, high control of ABMs leads to a low degree of external validity. To solve this problem, Betz presented a natural-language agent-based model of argumentation to simulate the argumentative opinion dynamics (Betz (2021)). The explicit stance used in traditional simulation of opinion dynamics can be represented by implicit argument from natural-language ABMAs.
2.3 Large Language Model
LLM such as GPT has been tested for multiple natural language processing tasks, and proved to be effective without gradient updates or fine-tuning (Brown et al. (2020)). Since GPT-3 was trained on large web corpora which include social media behavior, GPT-3 can predict user’s future behaviours, responses, or action plans with some accuracy. GPT was able to give human-like response to replace the expensive and time-consuming user studies (Sekulić et al. (2022)). Although the generated synthetic data was proved to exhibit less language variability, it may lead to better classification results when few data is available and resources are limited (Meyer et al. (2022)).
Recent study has use Large Language Models to simulate the online social interactions. LLMs are used to predict the popularity on the online social media (Park et al. (2022)), the distribution of votes for elections, the sentiment of news, and the voting of justices (Hamilton (2023)). In the work of Park et al. (2022), authors presented a social simulacra tool, SimReddit, which is a prototyping technique generating social behaviours using large language models. With the description of a community’s design, the social simulacra produce appropriate simulated behavior.
2.4 Reinforcement Learning
In 1898, reinforcement learning were laid by Thorndike (1898) through a cat behavior experiment employing a trial-and-error(TE) precedure. Originating from animal learning in psychology, RL has the capacity to mimic the human-like action selection. In 1989, Watkins & Dayan (1992) significantly advanced the field by integrating the theory of optimal control with temporal-difference(TD) learning, leading to the development of Q-learning. To address the curse of dimensionality in Q-learning, Mnih et al. (2015) combined deep learning with RL to create the deep Q-network(DQN), which notably surpassed professional human players in a range of 49 classic Atari games.
Recent advancements have extended RL’s application across various domains in the natural science, social sciences and engineering (Silver et al. (2016), Luo et al. (2017)). In the work of Sert et al. (2020), authors show that the combination of RL and ABM can create an artificial environment for policy makers to observe potential and existing behaviors associated to rules of interactions and rewards.
Figure 1: Simulated Environment Graph: ’Post’ is a small natural language message, formatting like Tweet. Participants within this environment, referred to as ’agents’, are interconnected through the ’follow’ relationship between each other. Evolution process shows the development of the whole network from $t$ to $t+1$. The process can be segmented into three part: Post Generation, Opinion Elicitation and Link Update.
3 Simulated Environment
3.1 Environment
In this section, we introduce the development of a simulated environment that replicates the dynamics of group discussions on online social media platforms. Discussions in this environment revolve around the topic, defined by posts that represent divergent viewpoints within the conversation.
The simulated environment system graph is shown in Figure 1. We will explain each part of the system step by step.
Within this simulated setting, a post is a small natural language message, formatting like Tweet. These posts are contributing to the whole discussion, each either supporting or opposing the central topic.
For instance, in our experiment, we initiate discussions using the topic:’Society with no gender’. Posts such as ’This incorrectly supposes that some current biological sex traits aren’t functional. I think a society without gender is a really bad idea.’ are categorized as ’con’ post to the topic. Conversely, posts like ’The absence of gender constructions would enable people to enjoy a better overall quality of life.’ support the topic, as a ’pro’ post.
Participants within this environment, referred to as ’agents’, are interconnected through the ’follow’ relationship between each other. Follow relationship is unidirectional. Each agent generates a post $\text{POST}^{t}_{i}$ at time t, influenced by their existing follow links and previous posts. Consequently, if $\text{agent}_{i}$ follows $\text{agent}_{j}$ at time $t$, then $\text{POST}^{t}_{j}$ potentially impacts $\text{POST}^{t+1}_{i}$.
At any given $t$, $\text{POST}^{t}_{i}$ fully determines the agent $i$’s opinion $\text{OPIN}^{t}_{i}$. The network of follow links between agents is then dynamically updated in response to these evolving $\text{OPIN}^{t+1}$.
The process can be segmented into three part: Post Generation, Opinion Elicitation and Link Update, as outlined in the following Algorithm1.
Algorithm 1 Simulated environment algorithm
for $t\in[1\ldots t_{\text{max}}]$ do
     for $i\in\text{AGENTS}$ do
         locate the follow link of agent $i$
         generate the post of agent $i$ ($\rightarrow\text{POST}_{i}^{t}$);
         elicit the opinion of agent $i$ ($\rightarrow\text{OPIN}_{i}^{t}$);
     end for
     update the follow link of AGENTS;
end for
3.2 Post Generation
This section explores the mechanics of large language models in content generation technology. These models predict the next-word in a sequence with a probabilistic approach given the prompt input.
The objective of the post generation stage is to generate the possible post of agent at the given time. This process incorporates the agent’s previous posts, and those from their followees.
The generated post must be concise a short sentence formatting like Tweet. Therefore we use the natural language description to describe the follow relationship, choose the format of web api including a designated stop sign as (’}’) to ensure the output aligns with the expected structure.
Consider the following example: if $\text{agent}_{i}$ follows $\text{agent}_{j}$, the prompt for generating $\text{POST}^{t+1}_{i}$ would be:
$\text{agent}_{i}$ posted tweets about ’a society without gender’: $\text{POST}^{t}_{i}$
$\text{agent}_{i}$ saw the following tweets about ’a society without gender’ on homepage{’data’:{’user’: $\text{agent}_{j}$,’text’: $\text{POST}^{t}_{j}$}. "
$\text{agent}_{i}$ shared the following tweet on this topic:{’data: {’user’: $\text{agent}_{i}$,’text’:’"
The first paragraph of the prompt describe the topic and refer the previous posts from $\text{agent}_{i}$. The second paragraph then suggests $\text{agent}_{i}$ follows $\text{agent}_{j}$, so the post of $\text{agent}_{j}$ will potentially impact the next post of $\text{agent}_{i}$. If $\text{agent}_{i}$ follows multiple agents in the group, the prompt will include more content that $\text{agent}_{i}$ can read from the homepage. The third paragraph ensure $\text{agent}_{i}$ still discuss the same topic and generation of tweet-like post. The name of agents, allowing for the observation of interaction between them, such as agreements or disagreements on the topic. An example of a generated post, demonstrating such an interaction, is as follow:
"I agree with both Grace Lee and Chloe Kim. Gender is an important aspect of our identity and should not be disregarded. However, everyone should also have the freedom to express themselves and live authentically. Instead of erasing gender, we should focus on promoting equality and challenging harmful gender stereotypes. Let’s create a society where individuals are free to express their gender identity without fear of discrimination or prejudice. #GenderEquality #BeYourself"
This response reflects incorporating with other agents and presenting a balanced perspective. It demonstrates that the generation of posts is not only topic-relevant but also reflective of the influences in a social media environment.
Moving to text generation process, key stages include ’encoding’, where the model assimilates and interprets the input text, identifying contextual patterns, and ’decoding’, where it generates new text from this analysis.
Crucial to the flexibility of these models in the decoding phase are key parameters such as ’temperature’ and ’top_p’. Temperature controls the randomness in the prediction, influencing creativity and coherence in the generated text. ’top_p’, controls the model to focus on a subset of the most probable next words, ensuring relevance and precision in the output. These two parameters balance the generated content between novelty and fidelity to the input context.
temperature top-p
Narrow 0.1 0.5
Creative 1.4 0.95
Table 1: GPT Decoding Setting
Table.1 shows the our setting for narrow case and creative case. Low temperature and low top-p lead to conservative, and narrow minded agents who’re sticking with the most obvious options when generating a post. The creative setting characterize agents who are much willing to take suprising turns.
3.3 Opinion Elicitation
The opinion elicitation process gives the probability that the agent expects a pro-claim rather than a con-claim given the perspective using perplexity calculation. The method we used is introduced in the work of Betz Betz (2021).
Perplexity effectively measures how well a probabilistic model predicts a sample, serving as an indicator of the model’s predictive capacity across the entire set of tokens within a corpus. In the context of natural language processing, perplexity is defined as the exponential of the average negative log-likelihood of a given sequence of tokens. Consider a tokenized sequence represented as $S=(w_{0},w_{1},...,w_{m})$. The perplexity of $S$ denoted as $\text{PPL}(S)$, is computed as:
$\text{PPL}(S)=\exp\left(-\frac{1}{m}\sum_{k=0}^{m}\log p_{\theta}(w_{k}|w_{<k})\right)$ (1)
In equation (1), $\log p_{\theta}(w_{k}|w_{<k})$ represents the logarithm of the likelihood of the $k$th token in the sequence, given all preceding tokens(denoted as $w_{<k}$) as per the probabilistic model in use.
When GPT processes a sentence, it calculates the log probability of each word given the previous words in the sentence which is $\log p_{\theta}(w_{k}|w_{<k})$ in the equation.
$\text{PPL}_{\text{CON}}(\text{POST}_{t}^{i})=\text{PPL}(S_{\text{CON}}|\text{POST}_{t}^{i}))$ (2)
Equation (2) shows how to compute ${\text{PPL}_{\text{CON}}(\text{POST}_{t}^{i})}$. Given a pre-defined determined sentence $S_{\text{CON}}$ like "society without gender is a really bad idea", the conditional perplexity shows the inverse probability of the output of the language model. The low perplexity result suggests that the language model predicts this particular sentence with a high level of certainty.
$\text{OPIN}_{t}^{i}=\frac{\text{PPL}_{\text{CON}}(\text{POST}_{t}^{i})}{\text{PPL}_{\text{PRO}}(\text{POST}_{t}^{i})+\text{PPL}_{\text{CON}}(\text{POST}_{t}^{i})}.$ (3)
In the equation (3) we give the definition of $\text{OPIN}_{t}^{i}$. The outcome is a float number in range $[0,1]$ which represent the polarity where 0 represents con and 1 represent pro. We will then categorise the number into five polarity categories: ’strong con’, ’con’, ’neutral’, ’pro’ and ’strong pro’.
3.4 Link Update
To establish the ‘Following’ relationship between users, we adopt the premise that individuals inclined to follow users who express similar opinions, as suggested by Zhou et al. (2009). Once we have the categorised polarity of each tweet, we then hypothesise that individuals with the same polarity category hold similar opinions.
Therefore, we explore two link methods: a Follow Dynamics and a Follow-Unfollow Dynamics. In both cases, if two users consistently align in the same opinion category over a few consecutive steps, they are considered likely to follow each other, with a follow probability. In Follow-Unfollow Dynamics, the exist follow link may also break with an unfollow probability, if the opinions of two people differ during the consecutive steps.
Figure 2: RL Loop. Reward: change of number of followers at each time step. State: 1)the current opinion states of the initial accounts followed by the target agent, 2)the number of followers the target agent currently has. Action: pre-defined five opinion polarity categories.
4 Reinforcement Learning
In the previous part, we have introduced the whole simulated environment. In this part, we detail the reinforcement learning setting employed for the model.
4.1 Model Design
We set a target agent in this group network without any initial followers. Then the target agent can actively choose current opinion state and post related posts. The objective is try to let the target agent get more followers within the determined time step.
The reinforcement learning model utilized in this framework is based on the q-learning algorithm, a widely used method in machine learning for environments characterized by uncertainty. This technique is used for developing optimal strategies through learning from interactions within such environments.
The RL Loop structure is shown in Figure 2. The target agent interacts with the previous introduced simulated environment. At each time step, the agent observes the state of the environment, selects an action, receives a reward from the environment, leading to a new environmental state.
The reward mechanism for this model is the change of number of followers at each time step, encouraging strategies that enhance follower acquisition.
The observed state comprises two main elements: 1)the current opinion states of the initial accounts followed by the target agent, 2)the number of followers the target agent currently has.
The action space available to the agents is pre-defined five opinion polarity categories: ’strong con’, ’con’, ’neutral’, ’pro’ and ’strong pro’. These categories correspond to actual post content, mapped through a dataset-generated dictionary. Upon choosing a category, the target agent randomly selects a tweet from that category.
4.2 Reinforcement Learning Algorithm
Algorithm 2 Reinforcement Learning Algorithm
Initialize Q-values $Q(s,a)$ arbitrarily for all state-action pairs $(s,a)$.
for each episode do
     initialize state $s$
     for $t\in[1\ldots t_{\text{max}}]$ do
         target agent choose action $a$ from state $s$ using policy derived from Q.
         for $i\in\text{AGENTS}$ do
              generate the post of agent $i$ ($\rightarrow\text{POST}_{i}^{t}$);
              elicit the opinion of agent $i$ ($\rightarrow\text{OPIN}_{i}^{t}$);
         end for
         update the follow link of AGENTS;
         observe reward $r$, and new state $s^{\prime}$.
         Update Q-value for $(s,a)$
         Update state $s\leftarrow s^{\prime}$.
     end for
end for
Referencing Algorithm 2, the process begins with the initialization of Q-values for all state-action pairs. In each episode, the state is initialized, and the agent embarks on a series of time steps. During each time step, the target agent selects an action from its current state, based on a policy derived from the Q-values. Alongside, each agent in the system generates posts and forms opinions. The algorithm then updates the follow links among agents and processes the resulting reward and new state. This leads to an update in the Q-value for the chosen action, and the agent’s state is updated accordingly.
The Q-value update is represented by the formula:
$Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$
In this formula, $\alpha$ represents the learning rate, $\gamma$ the discount factor, $r$ the reward, and $Q(s^{\prime},a^{\prime})$ the maximum estimated future reward. This update rule is key to the learning process, as it adjusts the Q-values based on the reward received and the potential future rewards, leading to more informed and strategic decision-making by the agent in subsequent steps.
5 Experiment
5.1 Simulated Environment Initialisation
The initiation of our simulated environment begins with the creation of initial posts, for which we use the dataset released by Betz (2021). The dataset is crawled from debating platform kialo.com, comprises various posts structured around specific topics. For our simulation, we select posts related to the topic ’Gender’ and topic ’Drug’.
(a) Gender
(b) Drug
Figure 3: Distribution of polarity value.
Employing the opinion elicitation process, we analyze the datasets to determine the polarity distribution. As shown in Figure 3, the distributions of polarity value in both cases cluster around the middle range. This distribution is indicative of the varying degrees of support or opposition to the topic within this dataset.
Given the polarity distributions, we categorise the posts into distinct opinion groups, aiming for an even distribution across these categories. The categorisations is shown in table 4 and table 5.
5.2 Simulated Environment Analysis
To examine the opinion evolution process within our simulated environment, we run two 50-step simulations using both Narrow and Creative agent setting using Gender topic. These simulations aim to demonstrate how different agent settings can influence the dynamics of the discussion.
Figure 4: Opinion Evolution Visualization on Narrow setting and Creative setting
Figure 4 illustrates the dynamics of opinion evolution under the influence of two simulation configurations: the Narrow setting and the Creative setting. Table 1 provides the specific Temperature and Top-p settings used for the opinion dynamics simulation.
Opinion evolution visualizations show the trajectory of opinion polarity score for 20 agents over 50 steps on Narrow setting and Creative Setting. Initially, the agents’ opinions range from 0.3 to 0.8. In the Narrow setting, we observe that opinion density increases over time, showing a tendency toward opinion convergence. Conversely, in the Creative setting, the opinion density remains more dispersed and chaotic throughout the entire time scale, with less apparent convergence.
To provide a more detailed view of opinion dynamics, opinion distributions at Time 0, 25, and 50 are presented. In the Creative setting, despite the initial high degree of opinion diversity, after 25 steps, some grouping starts to occur, but full convergence is not reached even by step 50. The discussion remains dynamic, with polarity scores fluctuating in the noisy range between 0.4 and 0.7. On the other hand, in the Narrow setting, while there is significant opinion diversity in the first 25 steps, this diversity gradually decreases, and by step 50, opinions cluster in the 0.45 to 0.6 range, indicating a clearer sign of convergence.
In both scenarios, the distribution of the final polarity scores shows a slight skew towards higher values across the time steps. However, in the Narrow setting, the opinions form a more centralized distribution by step 50, indicating stronger clustering and a more definitive convergence of opinions, as compared to the Creative setting, which remains more dispersed throughout the simulation.
(a) Follow Dynamics, Part-Observable
(b) Follow Dynamics, Full-Observable
(c) Follow-Unfollow Dynamics, Part-Observable
(d) Follow-Unfollow Dynamics, Full-Observable
Figure 5: Learning Curves of RL in Gender case, with different language model settings, link update methods, and observability settings.
(a) Follow Dynamics, Part-Observable
(b) Follow Dynamics, Full-Observable
(c) Follow-Unfollow Dynamics, Part-Observable
(d) Follow-Unfollow Dynamics, Full-Observable
Figure 6: Learning Curves of RL in Drug case, with different language model settings, link update methods, and observability settings.
5.3 Reinforcement Learning Setting
In this section, we explore the application of RL within our simulated environment. The primary research question is: Can RL algorithms identify an optimal strategy that enables the target agent to maximize its follower count within a predetermined number of steps?
By testing different setting, we found that limiting the action space and incorporating self-observation are key factors for stable RL result.
By restricting the action space, we reduce the complexity of the decision-making process, enabling the RL agent to focus on a more manageable set of actions. This helps prevent the model from exploring irrelevant or suboptimal actions that could introduce noise and instability into the learning process. Additionally, a smaller, well-defined action space allows for faster convergence, as the agent can more efficiently learn the consequences of its actions within the environment.
Incorporating self-observation further enhances stability by allowing the agent to continuously assess its own state and progress. This self-awareness helps the agent make more informed decisions, as it can adjust its behavior based on how past actions have affected its follower count and other relevant metrics. Self-observation also promotes consistency, as the agent is better equipped to identify patterns in the environment and avoid erratic or random behaviors. Together, limiting the action space and incorporating self-observation contribute to generating more stable and reliable outcomes, ensuring that the RL agent can effectively optimize its strategy within the defined constraints of the simulation.
To investigate stable RL results, we conducted tests across eight scenarios.
  1. 1.
    Language Model Settings:
    • We employed two types of settings for the language models of agents, as delineated in Table 1: Narrow and Creative. These settings impact the nature of posts generated by the agents.
  2. 2.
    Link Update Methods:
    • Follow Dynamics: In this case, the follow rate is 0.8, and the unfollow rate is 0. This setting allows us to observe the growth of the target agent’s network under a scenario where new follow links are relatively likely to form.
    • Follow-Unfollow Dynamics: In this more dynamic setting, the follow rate is still 0.8, and the unfollow rate is set as 0.5. This case allowing existing follow links to break, thus adding a layer of complexity to the network dynamics.
  3. 3.
    Observability Settings:
    • Part-Observable Case: In this scenario, the target agent initially follows only one other agent.
    • Full-Observable Case: Contrasting the Part-Observable setting, in this case, the target agent begins by following every agent in the network.
For the reinforcement learning, we use Gender case and Drug case, set number of episodes 500, max of steps per episode as 5. Other parameters are shown in Appendix-Parameters.
5.4 Reinforcement Learning Result
This subsection presents the outcomes of our RL experiments conducted across the eight scenarios. Figure 5 illustrates the results for the Gender scenario and Figure 6 shows the results for the Drug scenario. In both figures, the colored lines represent the average reward of five repeated experiments, and the shaded areas indicate the variance in the results.
In all scenarios, we observe convergence in the reward learning curves, indicating that the RL algorithm successfully identified optimal solutions within the given settings. Both the Gender and Drug cases exhibit similar learning curve patterns under different conditions, highlighting consistent performance across varying contexts.
One notable observation is that all Follow-Unfollow Dynamics cases exhibited greater variance in their learning curves compared to the Follow Dynamics counterparts under same setting. This implies that more complex scenarios pose greater challenges in reaching optimal results.
Cases with Full-Observable settings demonstrates more stable learning curve than the part-observable ones. This suggests that target agent with complete network visibility are likely to learn strategies more effectively and with greater stability than those with limited visibility.
The Narrow and Creative settings, when applied to the Partially-Observable cases, presented distinct learning curves. The Narrow setting consistently outperformed the Creative setting, as evidenced by higher reward values. However, in the Full-Observable scenarios, the learning curves for both Narrow and Creative settings align closely. A possible explanation is that, in Part-Observable scenarios, the unpredictable nature of agents in the Creative setting, leading to lower performance compared to the Narrow setting. Conversely, in Full-Observable scenarios, the agents are better able to adapt to and anticipate the Creative setting’s variable patterns.
In summary, these results highlight the impact of observability and complexity in reinforcement learning environments, with Full-Observable settings and simpler dynamics facilitating more stable and efficient learning outcomes. The Narrow setting’s predictable nature resulted in higher performance than the unpredictable Creative setting. However, when observability was not a constraint, as in the Full-Observable scenarios, both settings showed similar performance. This suggests that with full visibility, agents can effectively adapt to even unpredictable dynamics.
6 Conclusion
In this section, we reflect on our contribution, future work, ethical consideration, and limitations.
We have introduced a novel application of RL within a simulated discussion environment to model individual behavior in the context of social media. Our findings show that limiting the action space and incorporating self-observation are critical factors for generating stable results in dynamic environments. These insights demonstrate the adaptability of the RL model to diverse conditions and its effectiveness in identifying optimal strategies. Additionally, the RL results offer a deeper understanding of both individual and collective behavior on digital communication platforms.
Our future work would be directed towards enhancing the realism and variability of our simulated social media environment. This may involve incorporating more sophisticated interaction rules to better mirror the real-world social media dynamics. Additionally, we plan to discover how different character profiles of agents perform and interact within the network. This exploration aims to provide a deeper understanding of the varied behaviors and their impacts on the overall dynamics of social media interactions.
The potential application would be providing valuable insights for digital marketing strategists and policymakers in the dynamics of online opinion formation and influence.
However, we acknowledge that one potential negative social impacts could be the misuse of the simulated environment in manipulating public opinion. If used unethically, the technology may craft highly persuasive and targeted misinformation campaigns on social media platforms. Conversely, this technology can also be instrumental in detecting malicious activities in social networks.
One limitation of our research is the inherent unpredictability of actual social media discussions. Social dynamics are complex and unpredictable, and our rule-based simulated environment cannot fully replicate. Although the pre-trained RL algorithm demonstrates promising performs in the simulated environment, it presents one version of the actual possible outcome. Despite this, our research provides valuable insights into the mechanics of social interaction and opinion dynamics in online groups.
[…]
7 Appendix A: GPT API Setting
Creative Narrow
max_tokens 2000 2000
temperature 1.4 0.1
top_p 0.95 0.5
n 1 1
stop "}" "}"
model "gpt-3.5-turbo" "gpt-3.5-turbo"
Table 2: GPT Creative Parameters
Table 2 outlines the parameters set for the GPT API, distinguishing between ’creative’ case and ’narrow’ case configurations. Within these settings, higher temperature and top-p values are indicative of a setup geared towards fostering creativity. For tasks related to post generation and opinion elicitation, we employ the pre-trained GPT-3.5 model developed by OpenAI OpenAI (2022).
8 Appendix B: Reinforcement Learning Setting Parameters
Parameters for Reinforcement Learning training are listed in table 3.
learning rate 0.01
discount rate 0.99
epsilon 0.005
exploration rate 0.01
Table 3: RL Training Parameters
8.1 Appendix C: Opinion Example
In this section, we will explore illustrative examples of ’strong pro’ and ’strong con’ concerning the concept of a ’society with no gender.’ These examples are meant to elucidate the polarized viewpoints on this topic.
A ’strong pro’ argument is presented as follows:
"Any violent behavior targeted at the traits now associated with gender would be illogical, because that would assume that gender stereotypes would persist without gender. Such an assumption would almost certainly also include the academically false assumption that gender is biological."
Conversely, a ’strong con’ perspective is showed in the following statement:
"It won’t disappear. The idea of Sexist Violence in the eyes of society might, but that’s only because what we call Gender-based violence is just that… Violence. What you would see in a drop in Gender based violence, there would be a spike in generalized Violence to take its place. Its only taking a tag of something and putting it into a generalized pool of everything else. It helps no one and Dis-benefits everybody involved. So a society without gender is a really bad idea."
8.2 Appendix D: Opinion Category Setting
Category Polarity score Number of posts
Strong pro 1 – 0.64 83
Pro 0.58–0.64 89
Neutral 0.51–0.58 128
Con 0.45–0.51 93
Strong con 0 – 0.45 93
Table 4: Gender opinion category setting
Category Polarity score Number of posts
Strong pro 1 – 0.55 101
Pro 0.48–0.55 128
Neutral 0.42–0.48 145
Con 0.37–0.42 102
Strong con 0 – 0.37 184
Table 5: Drug opinion category setting